# Bumblebee gravity: spherically-symmetric solutions away from the potential minimum

### v2

@@ Abstract
In this work, we study a vector model of spontaneous spacetime-symmetry breaking coupled to gravity: the bumblebee model. The primary focus is on static spherically symmetric solutions. Complementing previous work on black hole solutions, we study the effects on the solutions when the vector field does not lie at the minimum of its potential. We first investigate the flat spacetime limit, which can be viewed as a modified electrostatic model with a nonlinear interaction term. We study the stability of classical solutions generally and in the spherically-symmetric case. We also find that certain potentials, based on hypergeometric functions, yield a Hamiltonian bounded from below. With gravity, we solve for the spherically-symmetric metric and vector field, for a variety of choices of the potential energy functions, including ones beyond the quadratic potential like the hypergeometric potentials. Special case exact solutions are obtained showing Schwarszchild-Anti de Sitter and Reissner-Nordstrom spacetimes. We employ horizon and asymptotic analytical expansions along with numerical solutions to explore the general case with the vector field away from the potential minimum. We discover interesting features of these solutions including naked singularities, repulsive gravity, and rapidly varying gravitational field near the source. Finally, we discuss observational constraints on these spacetimes using the resulting orbital behavior.
@@ I Introduction
Since the 1990s, there has been substantial work on experimental and observational searches for hypothetical spacetime-symmetry breaking [19, 20], motivated by theoretical scenarios in which it could arise [21, 22, 23, 24, 25]. While no statistically-significant signal for a breakdown of Lorentz, CPT, or diffeomorphism symmetry has yet been found, the “parameter” space of possible effects remains large [26, 27, 28, 12, 29]. Studies of the effects of spacetime-symmetry breaking on black holes have been published in the last decades. [^1] The literature is extensive, examples include Refs. [30, 31, 32, 33, 34, 35].
@@
The plan of study is the following. In Sec. II, we review the background on the vector model to be studied. First, we explore the flat spacetime limit in Sec. III, with specific solutions presented or characterized for various cases of the potential $V$. In particular, we establish some generic Hamiltonian properties, such as lower boundedness in Sec. III.1. The full field equations coupled to gravity are the subject of Sec. IV and Sec. V. Section IV discusses some generalities including energy theorems, a simple cosmological solution, the Bekenstein “no-hair” argument, and special cases with exact solutions.
@@ III.1 Field equations, Hamiltonian, and potential energy
where $b^{2}$ is a positive constant with the dimensions of mass squared. This potential has a minimum when $X=\mp b^{2}$, thus yielding a nonzero value for $B_{\mu}$ which can be spacelike, timelike, or even lightlike. Note that the symmetry is broken when the vacuum (quantum) state $|0>$ has the expectation value as in (3). Previous results have identified the quadratic potential (12) to be unstable, this can be seen from the Hamiltonian when the constraints are satisfied - in this case it is unbounded from below [80, 81]. Other potentials have been countenanced, such as a quartic potential [56].
@@
To our knowledge, none of these hypergeometric potentials have been considered before in a classical analysis paper, and not with gravity couplings. One can confirm that for a class of hypergeometric functions, the Hamiltonian for such potentials is bounded from below. Note that the minimum of $V$ can in fact be negative, which is interesting for cosmology and black holes as discussed later. In figure 1 we plot the scaled dimensionless version of the potential portion of the Hamiltonian density from (11),
@@
Polynomial $V=\frac{1}{n}\lambda_{n}X^{n}$ No
Hypergeometric $V=g\Lambda^{4}[M(n,2,X/\Lambda^{2})-1]$ Yes, $n\geq 3$
@@ III.2.3 Stability argument and numerical solutions
Equation (30) shows the stability argument for the quadratic potential, Eq. (31) applies the same reasoning to the hypergeometric potential $V=g\Lambda^{4}[M(3,2,z)-1]$, and Eq. (32) applies to $V=g\Lambda^{4}[M(4,2,z)-1]$, where $z=X/\Lambda^{2}$. Here, $f_{uu}$ denotes the second derivative of $f$ with respect to $u$, $\frac{d^{2}f}{du^{2}}$. Note that we can apply an argument about the trend of $f$ as $u$ tends toward smaller values because of the symmetry of the equations in (29) under the exchange $u\rightarrow-u$.
@@ IV Coupled gravity-vector solutions, Part I
This is the arbitrary spacetime generalization of the potential portion of the Hamiltonian density (14). So we already know that for hypergeometric potentials this is bounded from below in flat spacetime, although not positive definite (this holds even when $\vec{B}\neq 0$). This boundedness is valid in a local frame in the neighborhood of a point and can be demonstrated for all points. However, this fails for the other potential choices, with the exception of the massive potential. So it appears that at least in some cases, the weak energy condition will fail.
@@ VI.2 Summary
In flat spacetime, Sec. III, we established some general results for the spherical symmetry case. One key result is the analysis of the Hamiltonian for a generic potential $V(X)$, where $X=B_{\alpha}B^{\alpha}$, Eq. (11). We showed that certain hypergeometric potentials (13) have a lower bound on the Hamiltonian density, in contrast to polynomial potentials, as shown in Fig. 1.

### v3

@@ Abstract
In this work, we study a vector model of spontaneous spacetime-symmetry breaking coupled to gravity: the bumblebee model. The primary focus is on static spherically symmetric solutions. Complementing previous work on black hole solutions, we study the effects on the solutions when the vector field does not lie at the minimum of its potential. We first investigate the flat spacetime limit, which can be viewed as a modified electrostatic model with a nonlinear interaction term. We study the stability of classical solutions generally and in the spherically-symmetric case. We also find that certain potentials, based on hypergeometric functions, yield a Hamiltonian bounded from below for the case of fixed spatial vector magnitude. With gravity, we solve for the spherically-symmetric metric and vector field, for a variety of choices of the potential energy functions, including ones beyond the quadratic potential like the hypergeometric potentials. Special case exact solutions are obtained showing Schwarszchild-Anti de Sitter and Reissner-Nordstrom spacetimes. We employ horizon and asymptotic analytical expansions along with numerical solutions to explore the general case with the vector field away from the potential minimum. We discover interesting features of these solutions including naked singularities, repulsive gravity, and rapidly varying gravitational field near the source. Finally, we discuss observational constraints on these spacetimes using the resulting orbital behavior.
@@ I Introduction
Since the 1990s, there has been substantial work on experimental and observational searches for hypothetical spacetime-symmetry breaking [19, 20], motivated by theoretical scenarios in which it could arise [21, 22, 23, 24, 25]. While no statistically-significant signal for a breakdown of Lorentz, CPT, or diffeomorphism symmetry has yet been found, the parameter space of possible effects remains large [26, 27, 28, 12, 29]. Studies of the effects of spacetime-symmetry breaking on black holes have been published in the last decades. [^1] The literature is extensive, examples include Refs. [30, 31, 32, 33, 34, 35].
@@
The plan of study is the following. In Sec. II, we review the background on the vector model to be studied. First, we explore the flat spacetime limit in Sec. III, with specific solutions presented or characterized for various cases of the potential $V$. In particular, we establish some generic Hamiltonian properties in Sec. III.1. The full field equations coupled to gravity are the subject of Sec. IV and Sec. V. Section IV discusses some generalities including energy theorems, a simple cosmological solution, the Bekenstein “no-hair” argument, and special cases with exact solutions.
@@ III.1 Field equations, Hamiltonian, and potential energy
where $b^{2}$ is a positive constant with the dimensions of mass squared. This potential has a minimum when $X=\mp b^{2}$, thus yielding a nonzero value for $B_{\mu}$ which can be spacelike, timelike, or even lightlike. Note that the symmetry is broken when the vacuum (quantum) state $|0>$ has the expectation value as in (3). Previous results have identified the quadratic potential (12) to be unstable, this can be seen from the Hamiltonian when the constraints are satisfied - in this case it is unbounded from below [80, 81]. Other potentials have also been countenanced, such as a quartic potential [56].
@@
To our knowledge, none of these hypergeometric potentials have been considered before in a classical analysis paper, and not with gravity couplings. One can confirm that for a class of hypergeometric functions, the Hamiltonian for such potentials is bounded from below when the vector piece $\vec{B}||$ or $B_{0}$ is fixed. Note that the minimum of $V$ can in fact be negative, which is interesting for cosmology and black holes as discussed later. In figure 1 we plot the scaled dimensionless version of the potential portion of the Hamiltonian density from (11),
@@
Polynomial $V=\frac{1}{n}\lambda_{n}X^{n}$ No, $n$ even
Hypergeometric $V=g\Lambda^{4}[M(n,2,X/\Lambda^{2})-1]$ No, $n\geq 3$,
Yes, fixed $\vec{B}||$
@@
Despite this analysis, there remains a subtle aspect overlooked in a previous version of this manuscript. [^4] We thank Robertus Potting for pointing this out. This occurs along hyperbolas defined by constant $t=u^{2}-f^{2}$. In this case the hamiltonian is proportional to $e^{t}[1+f^{2}(3+t)+\tfrac{1}{2}t]-1$. From this expression for fixed t<3t<-3, the Hamiltonian can become arbitrarily negative for large $f$. Thus, the Hamiltonian for the hypergeometric function is not bounded from below in the generic $B_{\mu}$ case. Nonetheless, for subsets of the $B_{\mu}$ space, like constant $u$, it is bounded. Our focus in this work is on the latter.
A study of the local extrema of this function (16) with respect to $(f,u)$, shows that they exist at the points $f=\pm\sqrt{9\pm\sqrt{57}}/2,u=0$ and $f=0,u=0$. The value of ${\cal H}_{V}/(g\Lambda^{4})$ at these extrema is $0.235$ when $f=\pm\sqrt{9-\sqrt{57}}/2$, 1.092-1.092 when $f=\pm\sqrt{9+\sqrt{57}}/2$, and 0 when $f=0=u$. These extrema, which are all timelike values of $B_{\mu}$ turn out to be saddle points except for the solution $B_{\mu}=0$, which is a local minimum.
This analysis can be carried out for the $n>3$ potentials as well. We have checked for the same properties for the $n=4$, $n=5$ hypergeometric potentials, but a thorough study along these lines of the generic hypergeometric potentials remains open. In particular, conclusions for any $n$ may be drawn using the asymptotic properties of the hypergeometric functions [93].
@@ III.2.3 Stability argument and numerical solutions
Equation (30) shows the stability argument for the quadratic potential, Eq. (31) applies the same reasoning to the hypergeometric potential $V=g\Lambda^{4}[M(3,2,z)-1]$, and Eq. (32) applies to $V=g\Lambda^{4}[M(4,2,z)-1]$, where $z=X/\Lambda^{2}$. Here, $f_{uu}$ denotes the second derivative of $f$ with respect to $u$, $\frac{d^{2}f}{du^{2}}$.
@@ IV Coupled gravity-vector solutions, Part I
This is the arbitrary spacetime generalization of the potential portion of the Hamiltonian density (14). So we know that for hypergeometric potentials this is bounded from below in flat spacetime, when $\vec{B}||$ is fixed, but it is unbounded in the general case. These statements are valid in a local frame in the neighborhood of a point but can be demonstrated for all points. Positivity of (36) fails for many other potential choices, with the exception of the massive potential. So it appears that in many cases, the weak energy condition is violated.
@@ VI.2 Summary
In flat spacetime, Sec. III, we established some general results for the spherical symmetry case. One key result is the analysis of the Hamiltonian for a generic potential $V(X)$, where $X=B_{\alpha}B^{\alpha}$, Eq. (11). We discussed properties of hypergeometric potentials (13) and contrasted them with polynomial potentials. For example, see the plot of the Hamiltonian density for the case $\vec{B}=0$, shown in Fig. 1.